# SceneForge: Enhancing 3D-text alignment with Structured Scene Compositions

### v1

@@ 5.1 Proportion of Multi-Object Samples
Figure 6: Effect of varying $\alpha$ on average zero‑shot top‑1 accuracy. Performance peaks at $\alpha=0.5$, while higher values degrade performance.
The hyperparameter $\alpha$ determines the fraction of multi-object samples in each training batch. We analyze its impact by varying $\alpha$ from 0 to 1 and evaluating the average zero-shot top-1 accuracy on the zero-shot datasets (fig.˜6). Increasing $\alpha$ initially improves performance with respect to the baselines, with accuracy peaking around $\alpha=0.5$. However, for larger $\alpha$, performance declines, likely due to an excessive focus on compositional relationships at the expense of single-object understanding.
@@
SceneForge (N=2) 53.9 87.6 64.5 48.2
Table 4: Different 3D composition methods on zero‑shot cls. Our method achieves the best results across all benchmarks.
@@ 5.2 Do other composition functions work?
Results reported in table˜4 demonstrate that our approach outperforms other methods consistently across benchmarks. PointCutMix-R surpasses Uni3D (No Composition) on some metrics but remains inferior to our method, particularly compared to our more complex compositions ($N=3$) variant. Conversely, PointCutMix-K and PointMixUp significantly underperform compared to Uni3D. This disparity arises from each method’s effectiveness in preserving object semantics and spatial coherence. Our method constructs well-structured scenes aligning naturally with captions, whereas PointCutMix-R, despite retaining entire objects, mixes them randomly, generating overlapping and less organized scenes. PointCutMix-K and PointMixUp further disrupt semantic learning by fragmenting objects or interpolating features, resulting in unrealistic shapes poorly aligned with textual descriptions. These findings underscore the advantage of structured scene compositions, where maintaining clear object semantics and spatial relationships leads to enhanced generalization capabilities.
6 Limitations & Conclusion
We introduced SceneForge, a lightweight and model-agnostic framework that enhances 3D–text alignment through structured multi-object compositions. Our method improves contrastive learning across multiple 3D backbones—OpenShape, ViT-Lens, and Uni3D—without modifying architecture or supervision. Despite relying on simple synthetic arrangements, SceneForge yields significant improvements across a wide range of models and tasks, including zero-shot classification, segmentation, and spatial reasoning. Our findings show that even basic scene compositions enrich training and improve generalization. Looking forward, we believe that exploring more diverse and real-world inspired compositions—guided e.g., by object co-occurrence and spatial priors—offers a promising direction for advancing multimodal 3D understanding.

### v2

@@ 5.1 Proportion of Multi-Object Samples
Figure 6: Effect of varying $\alpha$ on average zero‑shot top‑1 accuracy.
The hyperparameter $\alpha$ determines the fraction of multi-object samples in each training batch. We analyze its impact by varying $\alpha$ from 0 to 1 and evaluating the average zero-shot top-1 accuracy on the zero-shot datasets (Figure˜6). Increasing $\alpha$ initially improves performance with respect to the baselines, with accuracy peaking around $\alpha=0.5$. However, for larger $\alpha$, performance declines, likely due to an excessive focus on compositional relationships at the expense of single-object understanding.
@@
SF-Uni3D (N=2) 53.9 87.6 64.5 48.2
Table 5: Different 3D composition methods on zero‑shot cls.
@@ 5.2 Do other composition functions work?
Results in Table˜5 show our method consistently outperforming all baselines. Although PointCutMix-R exceeds Uni3D on a most benchmarks, it still lags behind our method, even more considering the $N=3$ variant; PointCutMix-K and PointMixUp perform worse than Uni3D. This gap stems from how each approach handles object semantics and spatial coherence: our method builds structured scenes that align naturally with captions, whereas PointCutMix-R randomly mixes whole objects, creating overlaps, and PointCutMix-K and PointMixUp fragment or interpolate shapes, producing unrealistic, poorly described scenes. Overall, maintaining clear object semantics and spatial relationships through structured composition yields superior generalization.
5.3 Are simple relations enough?
To test if the learned spatial understanding generalizes beyond the simple pre-training relations, we isolate performance on ScanQA questions involving more complex, unseen spatial queries. Table 6 shows that our compositionally-trained models consistently outperform their baselines across all backbones and a variety of complex relations.
Table 6: Generalization to unseen spatial relations on ScanQA for all backbones.
Relation Type Metric OpenShape ViT-Lens Uni3D
Baseline SF $\Delta$ Baseline SF $\Delta$ Baseline SF $\Delta$
Attached To (21) CIDEr 54.5 61.5 +7.0 57.1 63.3 +6.2 57.9 66.6 +8.7
EM 14.1 17.1 +3.0 15.6 17.9 +2.3 16.5 20.8 +4.3
Sitting On (59) CIDEr 56.8 63.4 +6.6 59.0 65.1 +6.1 61.0 70.1 +9.1
EM 15.2 17.7 +2.5 16.6 18.4 +1.8 17.5 22.6 +5.1
Between (112) CIDEr 54.1 61.2 +7.1 56.8 62.9 +6.1 57.2 66.5 +9.3
EM 14.0 17.0 +3.0 15.5 17.9 +2.4 15.8 20.5 +4.7
Closest To (112) CIDEr 55.0 61.8 +6.8 57.5 63.5 +6.0 58.5 67.0 +8.5
EM 14.3 17.2 +2.9 15.8 18.0 +2.2 16.2 20.4 +4.2
In Front Of (246) CIDEr 56.1 62.5 +6.4 58.2 64.0 +5.8 60.3 68.3 +8.0
EM 14.9 17.6 +2.7 16.3 18.2 +1.9 17.1 21.8 +4.7
The comprehensive results in Table 6 show that SCENEFORGE provides consistent benefits across all three backbones. Each SF-variant significantly outperforms its respective baseline on all complex relation types. This indicates that our pre-training builds a robust spatial foundation that generalizes effectively to more nuanced relational queries, regardless of the underlying encoder architecture.
6 Limitations and Future Directions
While SceneForge consistently enhances 3D–text alignment across multiple backbones and tasks, we are aware of its limitations. First, our synthetic scene generation employs only three basic spatial relations, which, although diversified through LLM-based refinement, do not fully capture the complexity of natural environments. Future research could focus on more realistic and varied364compositions guided by learned object co-occurrence patterns and spatial priors. Second, due to computational constraints, we maintain a fixed 10 k-point budget for multi-object compositions, resulting in accuracy degradation for densely populated scenes (see Figure 3). Addressing this will require exploring larger point budgets or employing more sophisticated sampling techniques to preserve salient geometric features in complex scenarios. Third, although we leverage the lightweight Qwen2.5 model for refinement, the overhead is not always negligible and reducing it introduces a memory-time tradeoff. Finally, due to rendering costs, we could not pair each composition with a synthetic image, though extending the pipeline to incorporate rendered views for joint 2D–3D learning or studying alternative approaches, such as aligning 3D compositions to aggregations of the single image embeddings, remains a promising direction. Overall, addressing these limitations will significantly broaden the practical impact and robustness of compositional 3D–text learning methods.
Acknowledgments. This paper is supported by the PNRR-PE-AI FAIR project funded by the NextGeneration EU program. We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy) .